# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@ II.2.3 A numerical study of the accuracy: Hubbard model
Panels (b) and (c) in Fig. S3 show a comparison between the exact and SCI wavefunction amplitudes ordered by their magnitude for $U=1$ (metal) and $U=16$ (Mott insulator). For more concentrated wavefunctions ($U=16$) the SCI amplitudes are in better agreement with the exact amplitudes. For $U=1$ the number of relevant electronic configurations in the exact wavefunction greatly exceeds the $d=10^{3}$ configurations that SCI is allowed to use, explaining the poor performance of SCI in the prediction of the ground-state energy as shown in Fig. S3 (a).
II.3 Energy-variance extrapolation
It is expected that the accuracy of the HPC quantum estimator (with or without configuration recovery) will increase as the number of configurations used for the subspace expansion $d$ is increased. However, the convergence of the ground-state properties with $d$ is not expected to follow any specific functional relation. Therefore, attempting to analyze the convergence as a function of $d$ is not well motivated. Instead, the different eigenstate approximations obtained for different values of $d$ and different batches of samples $\mathcal{S}^{(k)}$ are used for an energy-variance extrapolation [48, 93, 94, 95, 96, 97]. Consider the approximate eigenstate $\left|\psi\right\rangle$, whose energy is given by $E=\langle\psi|\hat{H}|\psi\rangle$, and consider the exact eigenstate energy $E_{\textrm{T}}$. The difference between the approximate energy and $E_{\textrm{T}}$,
$\delta E=\langle\hat{H}\rangle-E_{\textrm{T}},$ (S58)
vanishes linearly with the Hamiltonian variance divided by the square of the variational energy [48]
$\frac{\Delta H}{E^{2}}=\frac{\langle\psi|\hat{H}^{2}|\psi\rangle-\langle\psi|\hat{H}|\psi\rangle^{2}}{E^{2}}.$ (S59)
In the main text and in this supplementary material we apply the energy-variance analysis to two different sets of energy-variance pairs. The first one is the energy-variance pairs obtained by the HCI procedure where different energies and variances are obtained by changing the cutoff parameter that indirectly controls the number of determinants in the subspace projection and diagonalization, resulting in different levels of accuracy. The second set of energy-variance pairs are those obtained from the HPC quantum estimator for different numbers of configurations $d$ as well as for different batches of sampled configurations $\mathcal{S}^{(k)}$.
III Experiments
This section begins by describing the quantum circuits that are employed to produce the configurations to which we apply the eigenstate solver. We describe how to set the circuit parameters from an efficient classical CCSD calculation, and show examples of the structure of the compiled circuits. We show numerical experiments comparing the performance of the circuits we use against the performance of a classically efficient reduction of the LUCJ circuits.
We also describe the setup for all of the experiments conducted in this work, highlighting the quantum and classical hardware used. We show the quantum device mappings, depth and number of gates in the circuits, and the number of measurement outcomes sampled. We also provide details about the classical computing resources.
III.1 Quantum circuits
In this work, we used the local-unitary cluster Jastrow (LUCJ) ansatz [40] to sample randomly distributed electronic configurations. LUCJ derives from the unitary cluster Jastrow (UCJ) ansatz, which has the form [41] of a product of $L$ layers,
$|\Psi\rangle=\prod_{\mu=1}^{L}e^{\hat{K}_{\mu}}e^{i\hat{J}_{\mu}}e^{-\hat{K}_{\mu}}|{\bf{x}}_{\mathrm{RHF}}\rangle\;,$ (S60)
@@ III.1 Quantum circuits
$\hat{K}_{\mu}=\sum_{pq,\sigma}K_{pq}^{\mu}\,\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{q\sigma}\;,\;\hat{J}_{\mu}=\sum_{pr,\sigma\tau}J_{p\sigma,q\tau}^{\mu}\,\hat{n}_{p\sigma}\hat{n}_{q\tau}\;.$ (S61)
In Eq. (S61), $p,q=0\dots N_{\mathrm{MO}}-1$ label molecular spatial orbitals and $\sigma,\tau$ label spin polarizations ($\alpha,\beta$ for spin-up and spin-down electrons respectively). $K_{pq}^{\mu}$/$J_{pq,\sigma\tau}^{\mu}$ has complex/real matrix elements and is anti-Hermitian/symmetric. The UCJ ansatz can be derived from a twice-factorized low-rank decomposition of the qUCCD ansatz [98, 41], and the $L$-product form is such that the exact full configuration interaction wavefunction can be obtained via Eq. (S60) [41, 39].
The local UCJ or LUCJ [40] introduces a “local” approximation of the UCJ ansatz, which makes the following modifications for opposite-spin and same-spin number-number terms:
$\begin{split}\sum_{pq}J_{p\alpha,q\beta}\hat{n}_{p\alpha}\hat{n}_{q\beta}&\rightarrow\sum_{p\in S}J_{p\alpha,p\beta}\hat{n}_{p\alpha}\hat{n}_{p\beta}\\ \sum_{pq}J_{p\sigma,q\sigma}\hat{n}_{p\sigma}\hat{n}_{q\sigma}&\rightarrow\sum_{pq\in S^{\prime}}J_{p\sigma,q\sigma}\hat{n}_{p\sigma}\hat{n}_{q\sigma},\end{split}$ (S62)
where $\sigma=\alpha,\beta$ and the sets $S,S^{\prime}$ are such that the quantum circuit implementing $e^{i\hat{J}_{\mu}}$ has depth $O(1)$ and only comprises $O(|S|+|S^{\prime}|)$ number-number “nn gates,” i.e. two-qubit unitaries of the form $U_{nn}(\varphi)=e^{-i\frac{\varphi}{4}(Z_{p}+Z_{q}-Z_{p}Z_{q})}$, acting on adjacent qubits $p,q$ in the topology of a certain device. For example, on a heavy-hex device, $S=\{4k\;,\;k=0\dots(N_{\mathrm{MO}}-1)/4\}$ and $S^{\prime}=\{(p,p+1)\;,\;p=0\dots N-2\}$ [40]. The circuits $e^{\pm\hat{K}_{\mu}}$, on the other hand, can be implemented by a Bogolyubov circuit acting on $N_{\mathrm{MO}}$ qubits and comprising $O(N_{\alpha}(N_{\mathrm{MO}}-N_{\alpha}))$ gates and depth $O(N_{\mathrm{MO}})$ [99, 100, 101]. Through the local approximation, the LUCJ Ansatz balances hardware friendliness and accuracy, the latter ultimately deriving from its connection to coupled cluster theory and adiabatic state preparation [40].
III.1.1 Initialization of LUCJ parameters
  1. 1.
  2. 2.
  3. 3.
    We extend the unitaries to the following matrices,
    $\tilde{U}_{y,pr}=\delta_{pa}\delta_{ri}U_{ai,y}\;,$ (S63)
    i.e. matrices where only the occupied/unoccupied block is non-zero.
  4. 4.
    We define the Hermitian operators
    $X_{\pm,y}=\frac{1\mp i}{2}\left(\tilde{U}_{y}\pm i\tilde{U}_{y}^{T}\right)$ (S64)
    and their eigenpairs
    $X_{\pm,y}V_{\pm,y}=g_{\pm,y}V_{\pm,y}\leavevmode\nobreak\ .$ (S65)
  5. 5.
    We define the operators
    $\begin{split}\left(J_{2y}\right)^{\sigma\tau}_{pr}&=\tau_{y}\left(g_{+,y}\right)_{p}\left(g_{+,y}\right)_{r}\\ \left(K_{2y}\right)^{\sigma}_{pr}&=\log(V_{+,y})_{pr}\\ \left(J_{2y+1}\right)^{\sigma\tau}_{pr}&=-\tau_{y}\left(g_{-,y}\right)_{p}\left(g_{-,y}\right)_{r}\\ \left(K_{2y+1}\right)^{\sigma}_{pr}&=\log(V_{-,y})_{pr}\leavevmode\nobreak\ .\\ \end{split}$ (S66)
  6. 6.
    We retain the first $L$ matrices $J$, $K$. This allows us to refine a non-local UCJ wavefunction [98, 41].
  7. 7.
    $\begin{split}&\left(J_{\mu}\right)^{\alpha\alpha}_{p,p+1}\;,\;p=0\dots N_{\mathrm{MO}}-2\\ &\left(J_{\mu}\right)^{\beta\beta}_{p,p+1}\;,\;p=0\dots N_{\mathrm{MO}}-2\\ &\left(J_{\mu}\right)^{\alpha\beta}_{p,p}\;,\;p=0\dots N_{\mathrm{MO}}-1\;,\;p\%4=0\leavevmode\nobreak\ .\\ \end{split}$ (S67)
III.1.2 Mapping to heavy-hex devices
The choice of retaining the elements $(J_{\mu})^{\alpha\beta}_{p,p}$ with $p=0\dots N_{\mathrm{MO}}-1\;,\;p\%4=0$ when implementing LUCJ on a heavy-hex device has an important technical motivation: on such devices, assuming a number of qubits greatly exceeding $N_{\mathrm{MO}}$, spin-up and spin-down orbitals can be mapped on two segments of adjacent qubits forming a “zig-zag” pattern and connected through an auxiliary qubit for $p=0,4,8,\dots$ as shown in the three rightmost panels of Fig. S9. Such a qubit layout allows implementing LUCJ with a minimal overhead of $\mathsf{SWAP}$ gates (two per auxiliary qubit and layer of LUCJ). However, on current devices with up to 133 qubits, for $N_{\mathrm{MO}}>21$ one cannot couple $N_{\mathrm{MO}}/4$ spin-orbitals with opposite spins through auxiliary qubits without incurring a significant overhead of $\mathsf{SWAP}$ gates, for the simple reason that a chain of 22 or more qubits is longer than the “diagonal” of the device. In such a situation, as shown in the rightmost panel of Fig. S9, the segments on which spin-up and spin-down qubits are mapped form two “tails” that are not connected by auxiliary qubits. This fact has two implications: (i) no more than 6 spin-orbitals with opposite spins can be coupled through auxiliary qubits, and (ii) if one retains the elements $(J_{\mu})^{\alpha\beta}_{p,p}$ with $p=0\dots N_{\mathrm{MO}}-1\;,\;p\%4=0$ and $p\leq 16$, the largest elements of $J_{\mu}$ may be discarded, yielding a lower-accuracy wavefunction. The first problem is a fundamental one, that can only be resolved with a substantially different mapping of fermionic degrees of freedom onto qubits and/or through the availability of larger devices. The second problem, on the other hand, has a simple solution, that we now describe. First, for any pair of spatial orbitals $p,r=0\dots N_{\mathrm{MO}}-1$, the orbital rotation
$\hat{S}_{pr}=e^{-i\frac{\pi}{2}\sum_{\sigma}\left(\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{r\sigma}+\hat{a}^{\dagger}_{r\sigma}\hat{a}^{\phantom{\dagger}}_{p\sigma}\right)}$ (S68)
$\hat{S}^{\dagger}_{pr}\left(\sum_{qs,\tau}M_{qs}\hat{a}^{\dagger}_{q\tau}\hat{a}^{\phantom{\dagger}}_{s\tau}\right)\hat{S}_{pr}=\sum_{qs,\tau}M^{\prime}_{qs}\hat{a}^{\dagger}_{q\tau}\hat{a}^{\phantom{\dagger}}_{s\tau}\;,\;M^{\prime}_{qs}=M_{S_{pr}(q),S_{pr}(s)}\;.$ (S69)
Three immediate implications of this fact are:
  • $\hat{S}^{\dagger}_{pr}\hat{n}_{q\tau}\hat{S}_{pr}=\hat{n}_{S_{pr}(q)\tau}$
    $\hat{S}^{\dagger}_{pr}\hat{J}_{\mu}\hat{S}_{pr}=\sum_{qs,\tau}(J^{\prime}_{\mu})^{\sigma\tau}_{q,s}\hat{n}_{q\sigma}\hat{n}_{s\tau}=\hat{J}^{\prime}_{\mu}\;,\;(J^{\prime}_{\mu})^{\sigma\tau}_{q,s}=({J}_{\mu})^{\sigma\tau}_{S_{pr}(q),S_{pr}(s)}\;.$ (S70)
$e^{-\hat{K}_{S}}\hat{J}_{\mu}e^{\hat{K}_{S}}=\hat{J}^{\prime}_{\mu}$ (S71)
$e^{\hat{K}_{\mu}}e^{i\hat{J}_{\mu}}e^{-\hat{K}_{\mu}}=e^{\hat{K}_{\mu}}e^{\hat{K}_{S}}e^{i\hat{J}^{\prime}_{\mu}}e^{-\hat{K}_{S}}e^{-\hat{K}_{\mu}}=e^{\hat{K}^{\prime}_{\mu}}e^{i\hat{J}^{\prime}_{\mu}}e^{-\hat{K}^{\prime}_{\mu}}$ (S72)
to obtain a UCJ operator with transformed orbital rotations $e^{-\hat{K}^{\prime}_{\mu}}$ and an opposite-spin density-density interaction whose largest elements in absolute value act on spatial orbitals $p=0,4,8,\dots,4\ell$. Sparsification of $(J^{\prime}_{\mu})^{\alpha\beta}_{q,s}$ then leads to retaining the $l$ dominant opposite-spin density-density interaction terms (as many as allowed by the size and topology of the available heavy-hex device) with a minimal overhead of $\mathrm{SWAP}$ gates.
III.1.3 Compiled circuits
@@
In Figures S4, S5,  S6, and  S7, we show the specific LUCJ circuits considered in this work. Each circuit is compiled into single-qubit
@@
III.1.4 Effect of the Jastrow terms in the accuracy
@@
In this section we study the impact of the two-body terms, $\exp(i\hat{J}_{\mu})$, of the LUCJ ansatz in the performance of the HPC quantum estimator. We compare the case where the two-body parameters are zero, ${J_{p\alpha,p\beta}=J_{p\alpha,p\alpha}=0}$, against the most general case where they can take arbitrary real values. Without the density-density interactions, the remaining circuit is a free-fermion evolution, i.e. the wavefunction amplitudes remain those of a single Slater determinant [82], so it can be efficiently simulated by a classical computer.
@@
We present a numerical study of that emulates the procedure without taking into account the effect of noise. Thus, there is no configuration recovery. For the quantum circuit simulations, we used the software library ffsim [42]. ffsim performs the state vector simulation within a subspace of fixed particle number and total $\hat{z}$ component of the spin, making the simulations much more efficient than a generic quantum circuit simulator.
@@
Three circuits are considered: LUCJ($L=1$), LUCJ($L=4$), and the LUCJ circuit without two-body terms, which we refer to as “Determinant”. We study the accuracy of the estimator in the dissociation of N2 (6-31G). For the LUCJ circuits, we consider those with heavy-hex-like connectivity in the density-density interactions; see Eq. (S67). The circuits are optimized to minimize the estimator energy, as described in Sec. II.1.3 using the COBYLA [69] optimizer. At each optimization step we consider $K=1$ batches of samples obtained from the set of sampled configurations $\mathcal{X}$, where $|\mathcal{X}|=10^{7}$. The value of $d$ at each point in the dissociation curve is chosen to coincide with the number of configurations resulting from a converged HCI calculation, carried out on PySCF, in the same system.
@@
Fig. S8 (a) shows the potential energy surface obtained from the HPC quantum estimator run on configurations generated by LUCJ($L=1$), LUCJ($L=4$) and the determinant circuits. Panel (b) in Fig. S8 shows the energy error from the three different circuits. For most bond lengths, the error of the HPC quantum estimator is largest for the determinant circuit, showing that the density-density interactions play a crucial role for the generation of relevant electronic configurations. This claim is also supported by the decrease of the error (on average) when increasing the number of LUCJ layers. We note that the optimization of the LUCJ circuits takes more iterations to converge than the optimization of the determinant circuits.
@@
Figure S8: Numerical and noiseless study of the accuracy of the SCI-based HPC quantum estimator obtained from different levels of approximation of the LUCJ circuit and samples obtained from an Slater determinant (limit when the sampling from the LUCJ circuit can be efficiently simulated). The system considered is the N2 dissociation in the 6-31G basis. The circuits are optimized to minimize the estimator energy. In the legend, $L$ refers to the number of layers of the LUCJ ansatz. (a) ground-state energy as a function of the bond length. (b) Difference between the exact diagonalization energy and the HPC quantum estimator energy as a function of the bond length.
III.2 Quantum and classical computational resources
III.2.1 Quantum experiment details

### v3

@@ 1.1.3 A numerical study of the accuracy: Hubbard model
Panels (b) and (c) in Fig. S1 show a comparison between the exact and SQD wavefunction amplitudes ordered by their magnitude for $U=1$ (metal) and $U=16$ (Mott insulator). For more concentrated wavefunctions ($U=16$) the SQD amplitudes are in better agreement with the exact amplitudes. For $U=1$ the number of relevant electronic configurations in the exact wavefunction greatly exceeds the $d=10^{3}$ configurations that SQD is allowed to use, explaining the poor performance of SCI in the prediction of the ground-state energy as shown in Fig. S1 (a).
1.2 An analytical lower bound to the probability of recovering families of configurations
  • $\varepsilon$; if $p$ is in the 1-sector of ${\bf{x}}_{r}$ and $x_{p}=1$.
  • $1-\varepsilon$; if $p$ is in the 1-sector of ${\bf{x}}_{r}$ and $x_{p}=0$.
  • $\varepsilon$; if $p$ is in the 0-sector of ${\bf{x}}_{r}$ and $x_{p}=0$.
  • $1-\varepsilon$; if $p$ is in the 0-sector of ${\bf{x}}_{r}$ and $x_{p}=1$.
We use these average values for our analytical derivations.
$\widetilde{\rho}=p^{L}\rho+\sigma~{},$ (S16)
  1. 1.
    $\frac{h\varepsilon}{(1-\varepsilon)b+\varepsilon(N-b+h)},$ (S21)
  2. 2.
$\begin{split}&P_{\textrm{recovery}}(M,N,b,\varepsilon)\geq\\ &\sum_{g=0}^{M-N-b}\sum_{h=0}^{N-b}\frac{\binom{M-N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(N-b+j)}\right)\\ &+\sum_{g=0}^{N-b}\sum_{h=0}^{M-N-b}\frac{\binom{N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(M-N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(M-N-b+j)}\right)\\ &\equiv F(M,N,b,\epsilon).\end{split}$ (S24)
Figure S2: Lower bound to the probability of recovering a particular bitstring with Hamming distance $2b$ from a reference approximation ${\bf{x}}_{r}$ bistring from a bistring sampled from the uniform distribution. The lower bound is shown as a function of the number of qubits. Different values of $b$ are shown, as indicated by the colorbar. The dashed black curve shows $1/2^{M}$, which is the probability of obtaining the desired bitsting directly from the uniform distribution. The circles connected by lines show the fraction of configurations for different families of bitstrings (as indicated by the colorbar) obtained by running the configuration recovery procedure on the samples from the quantum processor on the N2 molecule for the 6-31-G and cc-pVDZ bases sets. The inset shows the lower bound as a function of $b$ for $M=52$ qubits.
2 Optimization of the circuit parameters
The optimization of circuit parameters to produce better distributions in the space of electronic configurations is an avenue of improvement of the SQD framework. We develop a formalism for parameter optimization using the subspace energy as the cost function, in contrast to previous experiments that rely on standard VQE optimization frameworks  (?, ?). The HPC quantum estimator is constructed by running a diagonalization procedure in $K$ subspaces defined by different batches of sampled configurations, as described in the main text. The number of subspaces considered $K$ can go from $K=1$ to $K=K_{\textrm{max}}$, where $K_{\textrm{max}}=\binom{D}{d}$. Recall that $D$ is the dimensionality of the subspace of the Fock space spanned by configurations with the correct particle number and $d$ is the dimension of the SCI subspace. Note that $K_{\textrm{max}}$ is double-combinatorial in the number of spin-orbitals and electrons. Therefore, computing the HPC quantum estimator (with or without configuration recovery) for all possible sets of $d$ configurations is not efficient. Instead, we propose the use of a Monte Carlo estimator over batches of configurations $\mathcal{S}^{(k)}$ to optimize the energy of the HPC quantum estimator averaged over different sets of configurations.
where $E^{(k)}$ is defined in the main text, Eq. (15), as the SQD ground-state energy for the configurations in $\mathcal{S}^{(k)}$. An estimator to the double combinatorially-large summation in $\mathcal{E}(\theta)$ can be obtained by the Monte-Carlo unbiased estimator:
$\mathcal{E}(\theta)\approx\frac{1}{K}\sum_{k=1}^{K}E^{(k)}$ (S27)
whose Monte-Carlo estimator is given by:
Another practical issue with the gradient-based optimization of $\mathcal{E}(\theta)$ is the requirement to evaluate $\left\langle{\bf{x}}_{(i)}|\Psi_{\theta}\right\rangle$ and $\left\langle\partial_{\theta}\Psi_{\theta}|{\bf{x}}_{(i)}\right\rangle$, which rely in the Hadamard test for their evaluation. The implementation of the Hadamard test requires controlled versions of the circuits that realize the variational states. The implementation of such controlled unitaries requires circuit depths beyond the reach of current quantum processors.
An alternative approach is to optimize $\mathcal{E}(\theta)$ by explicit-gradient-free methods like COBYLA  (?) or simulated annealing. Gradient-free optimization of the circuit parameters is used in the numerical study of Sec. 3.0.2.
2.1 Orbital optimization
Orbital or single-particle basis rotations are a common approach to improve the accuracy of variational calculations of electronic structure systems. Orbital optimizations can be applied in conjunction with a wide variety of electronic structure methods, including complete active space diagonalizations  (?, ?, ?, ?), the density matrix renormalization group  (?, ?, ?), variational Monte Carlo  (?) techniques, and quantum computing variational approaches  (?, ?, ?). Moreover, orbital optimizations are common practice on SCI-based approaches as well  (?). More generally, orbital optimizations make variational approaches invariant under orbital rotations. For notation clarity we replace in this section the notation of the two-body integral $(pr|qs)$ by the abbreviation $h_{pqrs}$.
The implementation of the orbital optimizations in conjunction with SQD (with or without configuration recovery) follows the procedure presented in Ref.  (?). Orbital rotations consist of the application of the similarity transformation:
$\hat{\tilde{H}}=\hat{U}^{\dagger}(\kappa)\hat{H}\hat{U}(\kappa),$ (S32)
@@ 2.1 Orbital optimization
In this work, we choose the matrix that parametrizes the rotation to be real: $\kappa\in\mathbb{R}^{N_{\mathrm{MO}}\times N_{\mathrm{MO}}}$. To enforce unitarity in $\hat{U}(\kappa)$ we require that $\kappa_{pq}=-\kappa_{qp}$. According to Thouless’s Theorem  (?), the action of the similarity transformation defined by $\hat{U}(\kappa)$ transforms the creation operators according to:
using Einstein’s summation convention.
A variational procedure is used to search for the single-particle basis that yields the optimal description of the ground state, given the collection of all possible $K_{\textrm{max}}$ bare variational trial states $\left|\psi_{\Theta}^{(k)}\right\rangle$. The variational procedure is similar to the method presented in Sec. 2. In this setting, $\left|\psi_{\Theta}^{(k)}\right\rangle$ is the SCI ansatz given as the linear combination of the $d$ determinants in $\mathcal{S}^{(k)}$ (see Eq. (14)). The set of variational parameters $\Theta$ is composed of the circuit parameters $\theta$ defining $\left|\Psi_{\theta}\right\rangle$ together with the wavefunction amplitudes $c_{\bf{x}}^{(k)}$ in the subspace defined by the SCI configurations. Each variational state is dressed by the same single-particle orbital rotation
Gradient descent and its variants can be used to minimize both $\kappa$ and $\theta$ in $\mathcal{E}(\kappa,\Theta)$, while the optimization of $c_{{\bf{x}}}^{(k)}$ is carried out by the diagonalization procedure (see Sec. Eigenstate solver). Note that the sum in the average contains double-combinatorially many terms. Consequently, a Monte-Carlo based estimator is used to obtained an unbiased estimate (see Sec. 2). Gradients with respect to $\theta$ have the same expression as in Eq. (S29) replacing $\hat{H}$ by $\hat{\tilde{H}}$. Gradients with respect to the orbital rotations can be computed by the contraction of the bare one- and two-body reduced density matrices (1- and 2-RDMs) with the gradients of the one- and -two body integrals with respect to $\kappa_{pq}$ in Eq. (S38):
and the bare 1-RDMs and 2-RDMs are defined as
The gradients of the integrals with respect to the rotation parameters are computed using the automatic differentiation (AD) tools of the software package Jax  (?). Since the evaluation of the gradients with respect to $\kappa$ requires the evaluation of an intractable sum, we use an unbiased Monte-Carlo estimator for its evaluation:
See Sec. 4.3 to see the effect of orbital optimizations in the accuracy of SQD (with configuration recovery). The section shows results run on measurement outcomes obtained from the Heron processor, to study the dissociation of N2 at cc-pVDZ level of theory.
3 Additional information about experimental details
In this section we show examples of the structure of the compiled circuits. We show numerical experiments comparing the performance of the circuits we use against the performance of a classically efficient reduction of the LUCJ circuits.
We also describe the setup for all of the experiments conducted in this work, highlighting the quantum and classical hardware used. We show the quantum processor mappings, depth and number of gates in the circuits, and the number of measurement outcomes sampled. We also provide details about the classical computing resources.
3.0.1 Compiled circuits
@@
In Figures S3, S4,  S5, and  S6, we show the specific LUCJ circuits considered in this work. Each circuit is compiled into single-qubit
@@
3.0.2 Effect of the Jastrow terms in the accuracy
@@
In this section we study the impact of the two-body terms, $\exp(i\hat{J}_{\mu})$, of the LUCJ ansatz in the performance of SQD. We compare the case where the two-body parameters are zero, ${J_{p\alpha,p\beta}=J_{p\alpha,p\alpha}=0}$, against the most general case where they can take arbitrary real values. Without the density-density interactions, the remaining circuit is a free-fermion evolution, i.e. the wavefunction amplitudes remain those of a single Slater determinant  (?), so it can be efficiently simulated by a classical computer.
@@
We present a numerical study of that emulates the procedure without taking into account the effect of noise. Thus, there is no configuration recovery. For the quantum circuit simulations, we used the software library ffsim  (?). ffsim performs the state vector simulation within a subspace of fixed particle number and total $\hat{z}$ component of the spin, making the simulations much more efficient than a generic quantum circuit simulator.
@@
Three circuits are considered: LUCJ($L=1$), LUCJ($L=4$), and the LUCJ circuit without two-body terms, which we refer to as “Determinant”. We study the accuracy of the estimator in the dissociation of N2 (6-31G). For the LUCJ circuits, we consider those with heavy-hex-like connectivity in the density-density interactions; see Eq. (27). The circuits are optimized to minimize the estimator energy, as described in Sec. 2 using the COBYLA  (?) optimizer. At each optimization step we consider $K=1$ batches of samples obtained from the set of sampled configurations $\mathcal{X}$, where $|\mathcal{X}|=10^{7}$. The value of $d$ at each point in the dissociation curve is chosen to coincide with the number of configurations resulting from a converged HCI calculation, carried out on PySCF, in the same system.
@@
Fig. S7 (a) shows the potential energy surface obtained from SQD run on configurations generated by LUCJ($L=1$), LUCJ($L=4$) and the determinant circuits. Panel (b) in Fig. S7 shows the energy error from the three different circuits. For most bond lengths, the error of SQD is largest for the determinant circuit, showing that the density-density interactions play a crucial role for the generation of relevant electronic configurations. This claim is also supported by the decrease of the error (on average) when increasing the number of LUCJ layers. We note that the optimization of the LUCJ circuits takes more iterations to converge than the optimization of the determinant circuits.
@@
Figure S7: Numerical and noiseless study of the accuracy of the SCI-based HPC quantum estimator obtained from different levels of approximation of the LUCJ circuit and samples obtained from an Slater determinant (limit when the sampling from the LUCJ circuit can be efficiently simulated). The system considered is the N2 dissociation in the 6-31G basis. The circuits are optimized to minimize the estimator energy. In the legend, $L$ refers to the number of layers of the LUCJ ansatz. (a) ground-state energy as a function of the bond length. (b) Difference between the exact diagonalization energy and the SQD energy as a function of the bond length.
3.1 Quantum and classical computational resources
3.1.1 processor layouts